OpenStudy (anonymous):

Solve the equation 3(a + 3) – 6 = 21.

OpenStudy (anonymous):

it is 6

OpenStudy (anonymous):

3(a+3)-6=21 3(a+3)=21+6 3(a+3)=27 (a+3)=27/3 (a+3)=9 a=9-3 a=6

OpenStudy (anonymous):

Thanks that really helps.

OpenStudy (anonymous):

You're welcome.